OpenStudy (anonymous):

can anyone please help me on my homework. it only two question and i'm stuck.it Hard for me.

OpenStudy (anonymous):

\[1) 3,125\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

\[2)\left(\begin{matrix}125 \\ 8\end{matrix}\right)\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

Best reward for best answer.please show your working.

OpenStudy (mathmale):

Raince, it's important that you include the instructions with each problem you post. I admire you for using Equation Editor to post your expression, but without explanations it's hard to know what you mean. Is that (2/3) a factor? a multiplicand? an exponent? What does \[\left( \frac{ 125 }{ 8 } \right)\]

OpenStudy (mathmale):

mean to you? What does \[\left(\begin{matrix}125 \\ 8\end{matrix}\right) \]

OpenStudy (mathmale):

mean to you?

OpenStudy (anonymous):

you just have to find the value of the following.

OpenStudy (anonymous):

it means 25 over 8. squared with 2/3.'Rational expression'.

OpenStudy (anonymous):

2/3 is a exponent.

OpenStudy (mathmale):

\[\left(\begin{matrix}125 \\ 8\end{matrix}\right)\frac{ 2 }{ 3 } \] Now you tell me that you mean\[\left( \frac{ 125 }{ 89 } \right)^{2/3}.\] If you have any doubt regarding how to type in an expression in Equation Editor, use the Draw facility instead to avoid misinterpretations such as the one that hs happened here.

OpenStudy (mathmale):

\[125^{2/3}\] can be re-written as \[[125^{1/3}]^2\] Perhaps you recognize that 125 is the cube of 5. Therefore, 125 ^ (1/3) =5. Squaring the resulting expression results in 5^2 = 25. Hope this helps. Please re-write the other expression using the Draw facility, if at all possible. That will make it easier for someone else to help you. Unfortunately, I need to get off the Internet now. Best of luck to you. mathmale, Moderator

OpenStudy (anonymous):

:( thanks!

OpenStudy (anonymous):

i type it correctly.it 125/8 with the power of 2/3

OpenStudy (anonymous):

main topic is maipulating of exponent.

OpenStudy (skullpatrol):

$$\Huge a^m* a^n=a^{m+n}$$ $$\Huge (a^m)^n=a^{mn}$$ $$\Huge (ab)^m=a^m*b^m$$

OpenStudy (mathmale):

Your\[3,125\frac{ 4 }{ 5 }\] would be much clearer if typed as \[3,125^\frac{ 4 }{ 5 }\] My first move would be to determine whether or not 3125 is the 5th power of some number. \[3125^\frac{ 4 }{ 5 }=[3125^{1/5}]^4\] Try this:

OpenStudy (mathmale):

By all means, do the \[3125^{1/5} part first.\]